Errata for “Contraction-free sequent calculi for geometric theories with an application to Barr’s theorem”, Archive for Mathematical Logic (2003), vol. 42, pp. 389–401.

Page 391, 3^{rd} line of Section 3: Replace \overline{x} with \overline{z}. Page 391, 8^{th} line of Section 3: Replace

∀\overline{x}(P_{1}& … &P_{m} ⊃ ∃y_{1}M_{1} ∨ … ∨ ∃y_{n}M_{n})

with

∀\overline{z}(P_{1}& … &P_{m} ⊃ ∃x_{1}M_{1} ∨ … ∨ ∃x_{n}M_{n})

Page 392: Replace (in the last two lines of the inference tree)

⇒ P_{1}& … &P_{m} ⊃ ∃x_{1}M_{1} ∨ … ∨ ∃y_{n}M_{n}

with

⇒ P_{1}& … &P_{m} ⊃ ∃x_{1}M_{1} ∨ … ∨ ∃x_{n}M_{n}

and

⇒ ∀\overline{x}(P_{1}& … &P_{m} ⊃ ∃x_{1}M_{∨} … ∨ ∃x_{n}M_{n})

with

⇒ ∀\overline{z}(P_{1}& … &P_{m} ⊃ ∃x_{1}M_{∨} … ∨ ∃x_{n}M_{n})

Page 399, proof of Theorem 6: Add “rules for &, ∨,” to “Since B and C do not contain ⊃ or ∀, the derivation of A consists of geometric rules,”